# On a class of interdiction problems with partition matroids: complexity and polynomial-time algorithms

### v3

@@ 1.1 Motivation and background
Given two partition matroids, $X$ and $Y$, on the same weighted ground set $N$, we consider the following zero-sum game between a leader and a follower. In the first step, the leader selects an independent set $I$ of matroid $X$. In the second step, the follower selects an independent set $I^{\prime}$ of a matroid minor $Y\setminus I$, which is derived from $Y$ by excluding the elements in $I$. The follower’s objective is to select a set $I^{\prime}$ with a maximal possible weight, whereas the leader selects $I$ so as to minimize the weight of the follower’s independent set $I^{\prime}$. Hereafter, the aforementioned bilevel model is referred to as a partition matroid interdiction (PMI) problem. In addition to this general definition, we provide a standard bilevel integer linear programming (bilevel ILP) formulation of the PMI problem in Section 2.1.
@@
At the same time, the PMI problem leverages some concepts and ideas from network interdiction models with resource sharing [34, 35] and Stackelberg pricing models with multiple followers [5, 6]. For example, in the network interdiction models [34, 35] the leader has different types of resources and associated budget constraints, so that each arc in the network can be affected by multiple resources simultaneously. On the other hand, in Stackelberg pricing models [5, 6] the authors consider multiple followers, e.g., rational customers, who attempt to minimize their individual costs based on the leader’s prices. The relationship between our model and both knapsack interdiction and Stackelberg pricing problems are explored further in Sections 1.2-1.3.
@@ 1.2 Related literature
Finally, we refer to Shi et al. [40], who consider a specific class of bilevel mixed-integer linear programming (bilevel MILP) problems, where the follower’s decision variables are all binary. In this problem setting, it is assumed that the follower does not solve its optimization problem to global optimality, but seeks a locally optimal solution with respect to an $m$-flip neighborhood; see, e.g., [37]. Shi et al. [40] demonstrate that for fixed $m$ the resulting problem admits a single-level MILP reformulation of polynomial size. Furthermore, if the follower’s feasible set is a matroid, then any 2-flip locally optimal solution for the follower is proved to be globally optimal. Put differently, for this class of bilevel problems the proposed MILP reformulation with $m\geq 2$ yields an exact solution.
@@ 1.3 Our approach and contributions
As another line of research, in Section 3.3, we analyze a simple greedy algorithm for the leader. The algorithm starts with an empty independent set. At each step, it updates the leader’s blocking decision by adding a ground element that preserves independence and maximizes the respective change in the leader’s objective function value.
It is well known that the outlined greedy algorithm provides a constant-factor approximation for the problem of maximizing a monotone submodular function subject to cardinality or matroid constraints; see, e.g., [19, 36]. In contrast, we demonstrate that the PMI problem corresponds to minimizing a monotone submodular function subject to similar constraints. On a positive note, it is proved that, under a mild assumption, the greedy algorithm achieves an optimal solution for a particular version of the PMI problem where the follower’s feasible set is formed by a uniform matroid, i.e., $k_{f}=1$. However, by a counterexample, we show that when $k_{f}\geq 2$, the greedy algorithm fails to provide any constant-factor approximation for the optimal leader’s objective function value in the PMI problem.
@@ 2.1 Integer programming formulation
$\mathcal{I}_{Y}(\mathbf{x})=\Big{\{}\mathbf{y}\in\{0,1\}^{n}:\;\mathbf{y}\leq\mathbf{1}-\mathbf{x},\;\sum_{i\in F_{k}}y_{i}\leq f_{k}\quad\forall k\in K_{f}:=\{1,\ldots,k_{f}\}\Big{\}}.$ (2b)
@@
Parameter(s) Definition
$N:=\{1,\ldots,n\}$ a ground set
$\boldsymbol{\beta}\in\mathbb{R}_{+}^{n}$ a weight vector
$L_{k}\subset N$ and $l_{k}\in\mathbb{Z}_{+}$, $k\in K_{l}$ subsets of $N$ and budgets of the leader
$F_{k}\subset N$ and $f_{k}\in\mathbb{Z}_{+}$, $k\in K_{f}$ subsets of $N$ and budgets of the follower
Table 2: Summary of the notations used in formulation [PMI].
[…]
@@ 2.2 Proof of $NP$-hardness
Given an instance of [MIS-D], we construct an associated instance of the partition matroid interdiction problem [PMI] in the following way. Let $n=|V|\cdot|E|$, $k_{l}=|E|$, $k_{f}=|V|$ and assume that $l_{k}=1$ for $k\in\{1,\ldots,|E|\}$ and $f_{k}=1$ for $k\in\{1,\ldots|V|\}$. We define the leader’s and the follower’s feasible sets, respectively, as
@@
~$\widetilde{\mathcal{I}}_{Y}(\mathbf{x})=\Big{\{}\mathbf{y}\in\{0,1\}^{|V|\times|E|}:\;\mathbf{y}\leq 1-\mathbf{x},\;\sum_{e\in E}y_{v,e}\leq 1\quad\forall v\in V\Big{\}}.$ (3b)
@@
  • [PMI-D]: Given an integer $\widetilde{q}=|V|-q\in\mathbb{N}$, is there a blocking decision $\mathbf{x}\in\widetilde{\mathcal{I}}_{X}$ that satisfies
@@
Let $S\subseteq V$, $|S|\geq q$, be a set of vertices such that $r_{v}=0$ for each $v\in S$. We demonstrate that $S$ is an independent set of $G$, i.e., [MIS-D] admits a “yes” instance. Assume to the contrary that there exist two vertices, $v_{1}\in S$ and $v_{2}\in S$, that are incident to each other, i.e., there is an edge $e=(v_{1},v_{2})\in E$. By construction of $\boldsymbol{\widetilde{\beta}}$, we have $\widetilde{\beta}_{v_{1},e}=\widetilde{\beta}_{v_{2},e}=1$. However, since $\mathbf{x}\in\widetilde{\mathcal{I}}_{X}$, it can be observed that
$x_{v_{1},e}+x_{v_{2},e}\leq 1,$
and, hence, it cannot be the case that $\widetilde{\beta}_{v_{1},e}(1-x_{v_{1},e})=0$ and $\widetilde{\beta}_{v_{2},e}(1-x_{v_{2},e})=0$ simultaneously. As a result, either $r_{v_{1}}=1$ or $r_{v_{2}}=1$ that contradicts the definition of $S$.
@@ 3.1 Duality-based algorithm
$\sum_{k\in K_{f}}\max_{\mathbf{y}^{\mbox{\tiny(k)}}}\Big{\{}\boldsymbol{\beta}^{\mbox{\tiny(k)}\top}\mathbf{y}^{\mbox{\tiny(k)}}:\;\mathbf{0}\leq\mathbf{y}^{\mbox{\tiny(k)}}\leq\mathbf{1}-\mathbf{x}^{\mbox{\tiny(k)}},\;\mathbf{1}^{\top}\mathbf{y}^{\mbox{\tiny(k)}}\leq f_{k}\Big{\}}.$ (7)
@@
$\max_{\mathbf{y}^{\mbox{\tiny(k)}}}\sum_{i\in F_{k}}\beta_{i}y_{i}$ (8a)
@@
$\min_{\alpha_{k},\boldsymbol{\gamma}^{\mbox{\tiny(k)}}}\Big{\{}\alpha_{k}f_{k}\;+\;\sum_{i\in F_{k}}\gamma_{i}(1-x_{i})\Big{\}}$ (9a)
@@
$\begin{gathered}\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\;\sum_{k\in K_{f}}\;\min_{j_{k}\in\{0,\ldots,n\}}\Big{\{}\beta_{j_{k}}f_{k}\;+\;\sum_{i\in\widetilde{F}_{k}(j_{k})}(\beta_{i}-\beta_{j_{k}})(1-x_{i})\Big{\}}=\\ \min_{\mathbf{x}\in\,\mathcal{I}_{X}}\;\min_{\mathbf{j}\,\in S}\;\sum_{k\in K_{f}}\Big{\{}\beta_{j_{k}}f_{k}\;+\;\sum_{i\in\widetilde{F}_{k}(j_{k})}(\beta_{i}-\beta_{j_{k}})(1-x_{i})\Big{\}}=\\ \min_{\mathbf{j}\,\in S}\;\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\;\sum_{k\in K_{f}}\Big{\{}\beta_{j_{k}}f_{k}\;+\;\sum_{i\in\widetilde{F}_{k}(j_{k})}(\beta_{i}-\beta_{j_{k}})(1-x_{i})\Big{\}},\end{gathered}$ (13)
@@ 3.2 Dynamic programming-based algorithm
$\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\sum_{k\in K_{f}}\psi_{k}(\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}),$
@@
$\psi_{k}(\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}):=\max_{\mathbf{y}^{\mbox{\tiny(k}\mbox{\tiny)}}}\Big{\{}\boldsymbol{\beta}^{\mbox{\tiny(k}\mbox{\tiny)}\top}\mathbf{y}^{\mbox{\tiny(k}\mbox{\tiny)}}:\;\mathbf{0}\leq\mathbf{y}^{\mbox{\tiny(k}\mbox{\tiny)}}\leq\mathbf{1}-\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}},\;\mathbf{1}^{\top}\mathbf{y}^{\mbox{\tiny(k}\mbox{\tiny)}}\leq f_{k}\Big{\}},$
and the superscript “$(k)$” for $k\in K_{f}$ refers to the subset of indices $F_{k}$. Put differently, [PMI] can be divided into $k_{f}$ distinct stages, where at each stage $k\in K_{f}$ we need to identify an optimal blocking decision $\mathbf{x}^{*\mbox{\tiny(k}\mbox{\tiny)}}$ in the subset $F_{k}$ controlled by the follower.
@@
3 for $k=k_{f}$ and $\mathbf{r}_{k}\in R_{k}$
4 begin
5       $V_{k}(\mathbf{r}_{k})\longleftarrow\min_{\,\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}\in A^{\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})}\psi_{k}(\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}})$
6       $\mathbf{x}^{*\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})\in\operatorname{argmin}_{\,\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}\in A^{\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})}\psi_{k}(\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}})$
@@
10for $k\in\{k_{f}-1,\ldots,1\}$ and $\mathbf{r}_{k}\in R_{k}$
11 begin
12       $\mathbf{r}_{k+1}\longleftarrow\left(r_{k,k^{\prime}}-\sum_{i\in F_{k}\cap L_{k^{\prime}}}x_{i}\right)_{k^{\prime}\in K_{l}}$
13       $V_{k}(\mathbf{r}_{k})\longleftarrow\min_{\,\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}\in A^{\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})}\Big{\{}\psi_{k}(\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}})+V_{k+1}(\mathbf{r}_{k+1})\Big{\}}$
14       $\mathbf{x}^{*\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})\in\operatorname{argmin}_{\,\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}\in A^{\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})}\Big{\{}\psi_{k}(\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}})+V_{k+1}(\mathbf{r}_{k+1})\Big{\}}$
@@
19 $\mathbf{x}^{*\mbox{\tiny(1}\mbox{\tiny)}}\longleftarrow\mathbf{x}^{*\mbox{\tiny(1}\mbox{\tiny)}}(\mathbf{r}^{*}_{1})$

### v4

@@ 1.1 Motivation and background
Given two partition matroids, $X$ and $Y$, on the same weighted ground set $N$, we consider the following zero-sum game between a leader and a follower. In the first step, the leader selects an independent set $I$ of matroid $X$. In the second step, the follower selects an independent set $I^{\prime}$ from a matroid $Y\setminus I$, which is derived by excluding the elements in $I$ from both the ground set and the independent sets of $Y$. The follower’s objective is to select a set $I^{\prime}$ with a maximal possible weight, whereas the leader selects $I$ so as to minimize the weight of the follower’s independent set $I^{\prime}$. Hereafter, the aforementioned bilevel model is referred to as a partition matroid interdiction (PMI) problem. In addition to this general definition, we provide a standard bilevel integer linear programming (bilevel ILP) formulation of the PMI problem in Section 2.1.
@@
At the same time, the PMI problem leverages some concepts and ideas from network interdiction models with resource sharing [34, 35] and Stackelberg pricing models with multiple followers [5, 6]. For example, in the network interdiction models [34, 35] the leader has different types of resources and associated budget constraints, so that each arc in the network can be affected by multiple resources simultaneously. On the other hand, in Stackelberg pricing models [5, 6] the authors consider multiple followers, e.g., rational customers, who attempt to minimize their individual costs based on the leader’s prices. The relationship between our model and both knapsack interdiction and Stackelberg pricing problems is explored further in Sections 1.2-1.3.
@@ 1.2 Related literature
Finally, we refer to Shi et al. [42], who consider a specific class of bilevel mixed-integer linear programming (bilevel MILP) problems, where the follower’s decision variables are all binary. In this problem setting, it is assumed that the follower does not solve its optimization problem to global optimality, but seeks a locally optimal solution with respect to an $m$-flip neighborhood; see, e.g., [38]. Shi et al. [42] demonstrate that, for fixed $m$, the resulting problem admits a single-level MILP reformulation of polynomial size. Furthermore, if the follower’s feasible set is a matroid, then any 2-flip locally optimal solution for the follower is proved to be globally optimal. Put differently, for this class of bilevel problems the proposed MILP reformulation with $m\geq 2$ yields an exact solution.
@@ 1.3 Our approach and contributions
As another line of research, in Section 3.3, we analyze a simple greedy algorithm for the leader. The algorithm starts with an empty independent set. At each step, it updates the leader’s blocking decision by adding a ground element that preserves independence and maximizes the respective change in the leader’s objective function value. It is well known that the outlined greedy algorithm provides a constant-factor approximation for the problem of maximizing a monotone non-decreasing submodular function subject to a single capacity constraint or matroid constraints; see, e.g., [19, 37]. In contrast, we demonstrate that the PMI problem corresponds to minimizing a monotone non-increasing submodular function subject to similar constraints.
Submodularity and monotonicity offer a new perspective on the PMI problem. Thus, while minimizing a submodular function with a single capacity constraint is $NP$-hard in general [36], the PMI problem with a single capacity constraint for the leader, i.e., $k_{l}=1$, remains polynomially solvable due to the DP-based algorithm; recall Table 1. Moreover, given that the leader’s objective function is non-increasing, it is reasonable to analyze the greedy algorithm as a potential heuristic for the leader. Surprisingly, we demonstrate that, if the leader’s feasible set is an arbitrary matroid and the follower has a single capacity constraint, i.e., $k_{f}=1$, then the PMI problem can be solved exactly by the greedy algorithm. On the negative side, we present a counterexample showing that when $k_{f}\geq 2$, the greedy algorithm fails to provide any constant-factor approximation for the optimal leader’s objective function value.
@@ 2.1 Integer programming formulation
$\mathcal{I}_{Y}(\mathbf{x})=\Big{\{}\mathbf{y}\in\{0,1\}^{n}:\;\mathbf{y}\leq\mathbf{1}-\mathbf{x},\;\sum_{i\in F_{k}}y_{i}\leq f_{k}\quad\forall k\in K_{f}:=\{1,\ldots,k_{f}\}\Big{\}}.{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}$ (2b)
Parameter(s) Definition
$N:=\{1,\ldots,n\}$ a ground set
$\boldsymbol{\beta}\in\mathbb{R}_{+}^{n}$ a weight vector
$L_{k}\subset N$ and $l_{k}\in\mathbb{Z}_{+}$, $k\in K_{l}$ leader’s partition of $N$ and respective budgets of the leader
$F_{k}\subset N$ and $f_{k}\in\mathbb{Z}_{+}$, $k\in K_{f}$ follower’s partition of $N$ and respective budgets of the follower
Table 2: Summary of the notations used in formulation [PMI].
[…]
@@ 2.2 Proof of $NP$-hardness
Given an instance of [MIS-D], we construct an associated instance of the partition matroid interdiction problem [PMI] in the following way. Assume that $n=|V|\cdot|E|$, $K_{l}=\{1,\ldots,|E|\}$, $K_{f}=\{1,\ldots|V|\}$, $l_{k}=1$ for $k\in K_{l}$ and $f_{k}=1$ for $k\in K_{f}$. Also, let $V=\{v_{1},\ldots,v_{|V|}\}$, $E=\{e_{1},\ldots,e_{|E|}\}$ and
$L_{k}=\{(v,e_{k})\mid v\in V\},\,k\in K_{l},\,\mbox{ and }\,F_{k}=\{(v_{k},e)\mid e\in E\},\,k\in K_{f}.$
That is, the leader’s and the follower’s feasible sets, are defined respectively, as
@@
~$\widetilde{\mathcal{I}}_{Y}(\mathbf{x})=\Big{\{}\mathbf{y}\in\{0,1\}^{|V|\times|E|}:\;\mathbf{y}\leq 1-\mathbf{x},\;\sum_{e\in E}y_{v,e}\leq 1\quad\forall v\in V\Big{\}},$ (3b)
@@
  • [PMI-D]: Given an integer $\widetilde{q}=|V|-q\in\mathbb{Z}_{+}$, is there a blocking decision $\mathbf{x}\in\widetilde{\mathcal{I}}_{X}$ that satisfies
@@
Let $S\subseteq V$, $|S|\geq q$, be a set of vertices such that $r_{v}=0$ for each $v\in S$. We demonstrate that $S$ is an independent set of $G$, i.e., [MIS-D] admits a “yes” instance. Assume to the contrary that there exist two vertices, $v^{\prime}\in S$ and $v^{\prime\prime}\in S$, that are incident to each other, i.e., there is an edge $e=(v^{\prime},v^{\prime\prime})\in E$. By construction of $\boldsymbol{\widetilde{\beta}}$, we have $\widetilde{\beta}_{v^{\prime},e}=\widetilde{\beta}_{v^{\prime\prime},e}=1$. However, since $\mathbf{x}\in\widetilde{\mathcal{I}}_{X}$, it can be observed that
$x_{v^{\prime},e}+x_{v^{\prime\prime},e}\leq 1,$
and, hence, it cannot be the case that $\widetilde{\beta}_{v^{\prime},e}(1-x_{v^{\prime},e})=0$ and $\widetilde{\beta}_{v^{\prime\prime},e}(1-x_{v^{\prime\prime},e})=0$ simultaneously. As a result, either $r_{v^{\prime}}=1$ or $r_{v^{\prime\prime}}=1$ that contradicts the definition of $S$.
@@ 3.1 Duality-based algorithm
$\sum_{k\in K_{f}}\max_{\mathbf{y}^{\text{\tiny({k})}}}\Big{\{}\boldsymbol{\beta}^{\text{\tiny({k})}\top}\mathbf{y}^{\text{\tiny({k})}}:\;\mathbf{0}\leq\mathbf{y}^{\text{\tiny({k})}}\leq\mathbf{1}-\mathbf{x}^{\text{\tiny({k})}},\;\mathbf{1}^{\top}\mathbf{y}^{\text{\tiny({k})}}\leq f_{k}\Big{\}}.$ (7)
@@
$\max_{\mathbf{y}^{\text{\tiny({k})}}}\sum_{i\in F_{k}}\beta_{i}y_{i}$ (8a)
@@
$\min_{\alpha_{k},\boldsymbol{\gamma}^{\text{\tiny({k})}}}\Big{\{}\alpha_{k}f_{k}\;+\;\sum_{i\in F_{k}}\gamma_{i}(1-x_{i})\Big{\}}$ (9a)
@@
$\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\;\sum_{k\in K_{f}}\;\min_{j_{k}\in\{0,\ldots,n\}}\Big{\{}\beta_{j_{k}}f_{k}\;+\;\sum_{i\in\widetilde{F}_{k}(j_{k})}(\beta_{i}-\beta_{j_{k}})(1-x_{i})\Big{\}}=$
$\;\,\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\;\min_{\mathbf{j}\,\in S}\;\sum_{k\in K_{f}}\Big{\{}\beta_{j_{k}}f_{k}\;+\;\sum_{i\in\widetilde{F}_{k}(j_{k})}(\beta_{i}-\beta_{j_{k}})(1-x_{i})\Big{\}}=$ (13)
$\;\;\min_{\mathbf{j}\,\in S}\;\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\;\sum_{k\in K_{f}}\Big{\{}\beta_{j_{k}}f_{k}\;+\;\sum_{i\in\widetilde{F}_{k}(j_{k})}(\beta_{i}-\beta_{j_{k}})(1-x_{i})\Big{\}},$
@@ 3.2 Dynamic programming-based algorithm
$\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\sum_{k\in K_{f}}\psi_{k}(\mathbf{x}^{\text{\tiny({k})}}),$
@@
$\psi_{k}(\mathbf{x}^{\text{\tiny({k})}}):=\max_{\mathbf{y}^{\text{\tiny({k})}}}\Big{\{}\boldsymbol{\beta}^{\text{\tiny({k})}\top}\mathbf{y}^{\text{\tiny({k})}}:\;\mathbf{0}\leq\mathbf{y}^{\text{\tiny({k})}}\leq\mathbf{1}-\mathbf{x}^{\text{\tiny({k})}},\;\mathbf{1}^{\top}\mathbf{y}^{\text{\tiny({k})}}\leq f_{k}\Big{\}},$
and the superscript “$(k)$” for $k\in K_{f}$ refers to the subset of indices $F_{k}$. Put differently, [PMI] can be divided into $k_{f}$ distinct stages, where at each stage $k\in K_{f}$ we need to identify an optimal blocking decision $\mathbf{x}^{*\text{\tiny({k})}}$ in the subset $F_{k}$ controlled by the follower.
@@
3 $k\longleftarrow k_{f}$
4 for  $\mathbf{r}_{k}\in R_{k}$
5       $V_{k}(\mathbf{r}_{k})\longleftarrow\min_{\,\mathbf{x}^{\text{\tiny({k})}}\in A^{\text{\tiny({k})}}(\mathbf{r}_{k})}\psi_{k}(\mathbf{x}^{\text{\tiny({k})}})$
6       $\mathbf{x}^{*\text{\tiny({k})}}(\mathbf{r}_{k})\in\operatorname{argmin}_{\,\mathbf{x}^{\text{\tiny({k})}}\in A^{\text{\tiny({k})}}(\mathbf{r}_{k})}\psi_{k}(\mathbf{x}^{\text{\tiny({k})}})$
@@ 3 $k\longleftarrow k_{f}$
10for $k\in\{k_{f}-1,\ldots,1\}$
11       for  $\mathbf{r}_{k}\in R_{k}$
12             $\boldsymbol{\theta}_{k}(\mathbf{r}_{k},\mathbf{x}^{\text{\tiny({k})}})\longleftarrow\left(r_{k,k^{\prime}}-\sum_{i\in F_{k}\cap L_{k^{\prime}}}x_{i}\right)_{k^{\prime}\in K_{l}}$
13             $V_{k}(\mathbf{r}_{k})\longleftarrow\min_{\,\mathbf{x}^{\text{\tiny({k})}}}\Big{\{}\psi_{k}(\mathbf{x}^{\text{\tiny({k})}})+V_{k+1}\big{(}\boldsymbol{\theta}_{k}(\mathbf{r}_{k},\mathbf{x}^{\text{\tiny({k})}})\big{)}:\mbox{ }\mathbf{x}^{\text{\tiny({k})}}\in A^{\text{\tiny({k})}}(\mathbf{r}_{k})\Big{\}}$
14             $\mathbf{x}^{*\text{\tiny({k})}}(\mathbf{r}_{k})\in\arg\min_{\,\mathbf{x}^{\text{\tiny({k})}}}\Big{\{}\psi_{k}(\mathbf{x}^{\text{\tiny({k})}})+V_{k+1}\big{(}\boldsymbol{\theta}_{k}(\mathbf{r}_{k},\mathbf{x}^{\text{\tiny({k})}})\big{)}:\mbox{ }\mathbf{x}^{\text{\tiny({k})}}\in A^{\text{\tiny({k})}}(\mathbf{r}_{k})\Big{\}}$
@@
20 $\mathbf{x}^{\text{\tiny(1)}}\longleftarrow\mathbf{x}^{\text{\tiny(1)}}(\mathbf{r}^{*}_{1})$